OpenStudy (anonymous):

Rationalise 1/(1+2(3)^(1/2)-(2)^(1/2))

OpenStudy (anonymous):

What is the conjugate of (1+2(3)^(1/2)-(2)^(1/2))

jhonyy9 (jhonyy9):

what is 3^(1/2) = how can you rewrite it ?

jhonyy9 (jhonyy9):

you need rewrite it your exercise without exponents ok ?

jhonyy9 (jhonyy9):

so for example if x^(1/2) = sqrt(x) than how can you rewriting your exercise ?

OpenStudy (anonymous):

1/(1+2sqrt(3)-sqrt(2))